OpenStudy (anonymous):

Doug entered a canoe race. He rowed 3 1/2 miles in 1/2 hour. What is his average speed in miles per hour?

OpenStudy (therealmeeeee):

i will do my best too help u

OpenStudy (therealmeeeee):

i dont know this stuff srry i mean ikw it im not good at it but have this dude help u

OpenStudy (therealmeeeee):

@TuringTest

jigglypuff314 (jigglypuff314):

miles per hour = miles/hour so 3 1/2 miles in 1/2 hour --> (3 1/2)/(1/2) =

jigglypuff314 (jigglypuff314):

sorry, if (3 1/2)/(1/2) = is hard to read then: \[\frac{ 3.5 }{ 0.5 } =\]